# Online Fair Allocations with Binary Valuations and Beyond

### v1

@@ 17      $A_{i}=A_{i}\cup\{e\}$;
$\lambda=2$. For the arriving item $e_{2}$ when $\lambda=2$, let us consider four cases about $\mu$.
  • When $\mu=1$, item $e_{2}$ is directly allocated to agent $j$. For agent $i$, her bundle is EF1 since after removing $e_{2}$ from $A_{j}$, we have $v_{i}(A_{i})\geq v_{i}(A_{j}\cup\{e_{1}\})=v_{i}(A_{j})+a_{i}$. For agent $j$, her new bundle is EF since we have $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$, $v_{j}(e_{1})=b_{j}\geq v_{j}(g_{2})$ and $v_{j}(e_{2})\geq v_{j}(g_{1})=a_{j}$. Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated.
  • When $\mu=2$, we need to consider two subcases. If $v_{i}(e_{2})=b_{i}$, agent $i$ picks it. For agent $i$, her new bundle is EF since we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}$ and $v_{i}(e_{2})=b_{i}=v_{i}(e_{1})$. For agent $j$, her bundle is $\frac{1}{2}$-EF1 since we have $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\setminus\{g\})$ for some $g\in A_{j}$ and $v_{j}(e_{1})=b_{j}\geq v_{j}(e_{2})$. Thus, this allocation is $\frac{1}{2}$-EF1 and satisfies the properties in Observation 3. Additionally, we set $\lambda=1$ and continue executing this algorithm. If $v_{i}(e_{2})=a_{i}$, it is allocated to agent $j$. For agent $i$, her bundle is EF1 since after removing $e_{1}$ from $A_{j}$,we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}=v_{i}(A_{j}\cup\{e_{2}\})$. For agent $j$, her new bundle is trivially EF since we have $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$, $v_{j}(e_{1})=b_{j}\geq v_{j}(g_{2})$ and $v_{j}(e_{2})\geq a_{i}=v_{j}(g_{1})$. Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated.
  • When $\mu=3$, let us also consider two subcases. If $v_{j}(e_{2})=b_{j}$, agent $j$ picks it. For agent $i$, her bundle is EF1 since after removing $e_{2}$ from $A_{j}$, we have $v_{i}(A_{i})\geq v_{i}(A_{j}\cup\{e_{1}\})=v_{i}(A_{j})+a_{i}$. For agent $j$, her new bundle is EF since $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$, $v_{j}(e_{2})=b_{j}\geq v_{j}(g_{2})$ and $v_{j}(e_{1})=a_{j}=v_{j}(e_{2})$. Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated. If $v_{j}(e_{2})=a_{j}$, agent $i$ picks it. For agent $i$, her new bundle is EF since we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}$ and $v_{i}(e_{2})=a_{i}=v_{i}(e_{1})$. For agent $j$, her bundle is $\frac{1}{2}$-EF1 since we have $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\setminus\{g\})$ for some $g\in A_{i}$ and $v_{j}(e_{1})=a_{j}=v_{j}(e_{2})$. Thus, this allocation is $\frac{1}{2}$-EF1 and satisfies the properties in Observation 3. Additionally, we set $\lambda=1$ and continue executing this algorithm.
  • When $\mu=4$, it is directly allocated to agent $j$. For agent $i$, her bundle is EF since $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}$ and $v_{i}(e_{1})=b_{i}\geq v_{i}(e_{2})$. For agent $j$, her new bundle is $\frac{1}{2}$-EF1 since $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\setminus\{g\})$ for some $g\in A_{i}$ and $v_{j}(e_{2})\geq v_{j}(e_{1})=a_{j}$. Thus, this allocation is $\frac{1}{2}$-EF1 and satisfies the properties in Observation 3. Additionally, we set $\lambda=1$ and continue executing this algorithm.
@@
Proof of Theorem 5.
In the main body of Algorithm 3, we follow a similar way of offline envy graph procedures to allocate items. The precondition of Algorithm 3 is that there is no envy cycle between two agents. Initially, the empty allocation is EF1 without an envy cycle. When there is no envy between two agents, we allocate the arriving item to one agent according to different conditions. When there exists an envy between two agents, we have two cases. One is that if we allocate it to the unenvied agent, the envy cycle does not appear. We will follow this procedure. The other is that if we allocate it to the unenvied agent, the envy cycle will appear, and we terminate Algorithm 3. In the online setting, we cannot swap bundles to eliminate the envy cycle. Thus, to enable the Algorithm 3 to be executed normally, Algorithms 4 and 5 help us to manually break the possible envy cycle, where, by Lemmas 1 and 2, we guarantee that $\frac{1}{2}$-EF1 always holds. When the allocation is EF1 and there is no envy cycle, we continue executing Algorithm 3. By the induction, we can conclude that the output allocation of Algorithm 3 is always EF1 without an envy cycle. Combining Lemmas 1, 2, and the description of Algorithm 3, it can be seen that we can guarantee an NW allocation that is $\frac{1}{2}$-EF1. As for MMS, by the implication between EF1 and MMS [21], i.e., $\alpha$-EF1 $\Rightarrow$ $\frac{\alpha}{(n-1)\alpha+1}$-MMS, where $n$ is the number of agents, it is clear that the output allocation is $\frac{\frac{1}{2}}{\frac{1}{2}+1}=\frac{1}{3}$-MMS. ∎
Input: An instance $(T,N,(v_{1},v_{2}))$ with additive bi-valued valuation functions
Output: An non-wasteful approximate EF1 and approximate MMS allocation $A$
@@
2Initialize: $\lambda=1$ and $\mu=0$;
3 when item $e\in T$ arrives do
4
5      if $\lambda=1$ then
6
@@ 9                  $A_{j}=A_{j}\cup\{e\}$;
10                  Terminate this algorithm and execute the adapted envy-graph procedures (Algorithm 3);
@@
12
13                  if $v_{j}(e)=b_{j}$ then
14
15                        $A_{j}=A_{j}\cup\{e\}$ and $\mu=1$;
@@ 31                        $A_{j}=A_{j}\cup\{e\}$;
32                  else
33
34                        $A_{i}=A_{i}\cup\{e\}$;
35                        Terminate this algorithm and execute the deep-breaking-cycle algorithm (Algorithm 5);
@@
39
40            Terminate this algorithm and execute the adapted envy-graph procedures (Algorithm 3);
41       end if
42
43
return $A$;
Algorithm 4 Preliminary-Breaking-Cycle Algorithm
Input: An instance $(T,N,(v_{1},v_{2}))$ with additive bi-valued valuation functions
Output: An non-wasteful approximate EF1 and approximate MMS allocation $A$
1 Initialize: $\lambda=1$ and $\mu=0$;
2 when item $e\in T$ arrives do
3
4      if $\lambda=1$ then
5
6            if $v_{i}(e)=a_{i}$ and $v_{j}(e)=b_{j}$ then
7                   $A_{j}=A_{j}\cup\{e\}$ and $\mu=1$;
8             else if $v_{i}(e)=b_{i}$ and $v_{j}(e)=b_{j}$ then
9
10                  $A_{j}=A_{j}\cup\{e\}$ and $\mu=2$;
11             else if  $v_{i}(e)=a_{i}$ and $v_{j}(e)=a_{j}$  then
12                   $A_{j}=A_{j}\cup\{e\}$ and $\mu=3$;
13             else
14                   $A_{i}=A_{i}\cup\{e\}$ and $\mu=4$;
15             end if
16            $\lambda=2$;
17      else
18             if $\mu=1$ then
19                   $A_{j}=A_{j}\cup\{e\}$;
20             else if  $\mu=2$  then
21                   if $v_{i}(e)=b_{i}$ then
22
23                        $A_{i}=A_{i}\cup\{e\}$;
24                        $\lambda=1$;
25                  else
26                         $A_{j}=A_{j}\cup\{e\}$;
27                   end if
28
29             else if  $\mu=3$  then
30                   if $v_{j}(e)=b_{j}$ then
31
32                        $A_{j}=A_{j}\cup\{e\}$;
33                  else
34
35                        $A_{i}=A_{i}\cup\{e\}$;
36                        $\lambda=1$;
37                   end if
38
39             else
40
41                  $A_{j}=A_{j}\cup\{e\}$;
42                  $\lambda=1$;
43             end if
44
45            if $\lambda=2$ then
46                   Terminate this algorithm and execute the adapted envy-graph procedures (Algorithm 3);
47
48       end if
49
50
51
52return $A$;
Algorithm 5 Deep-Breaking-Cycle Algorithm
Then, we present the impossible results to show that our approximation guarantees for EF1 and MMS are tight, and there is no compatibility between EF1 or MMS, and USW.
@@ 46                   Terminate this algorithm and execute the adapted envy-graph procedures (Algorithm 3);
For two agents with additive bi-valued valuation functions, no deterministic online algorithm can compute an NW allocation that is $(\frac{1}{2}+\epsilon)$-EF1 or $(\frac{1}{3}+\epsilon)$-MMS for any $\epsilon>0$.
@@
$t=1$ $t=2$ $t=3$ $t=4$
agent $1$ $\epsilon$ $1$ $\epsilon$ 1
@@
$t=1$ $t=2$ $t=3$ $t=4$ $t=5$
agent $1$ $\epsilon$ $1$ $\epsilon$ $\epsilon$ $1$
agent 2 $\epsilon$ $\epsilon$ $\epsilon$ 1 $1$
@@
$t=1$ $t=2$ $t=3$ $t=4$
agent $1$ $\epsilon$ $1$ $\epsilon$ $\epsilon$
@@
Table 4: The impossibility results of EF1 and MMS for additive bi-valued valuations for two agents.
Consider an online allocation instance with two agents $\{1,2\}$. For each agent $i\in\{1,2\}$, $v_{i}(e)\in\{\epsilon,1\}$, where $10<\epsilon<$ 1. The value of each arrived item is shown in Table 4. When $t=1$, the item with the value of $\epsilon$ for both agents arrives. Without loss of generality, we allocate $e_{1}$ to agent 1. When $t=2$, the arriving item has the value of 1 for agent 1 and $\epsilon$ for agent 2. To avoid 0-EF1 and 0-MMS, it should be allocated to agent 2. Then, when $t=3$, the arriving item has the value of $\epsilon$ for both agents. Let us consider two different choices of this item.
@@
For the deterministic allocation of indivisible goods, no online algorithm can compute an NW allocation that is $\alpha$-EF1 or $\alpha$-MMS, and $\beta$-max-USW for any $\alpha,\beta>0$, even for two agents with additive bi-valued valuation functions.
@@
Consider an online allocation instance with two agents $\{1,2\}$. Assume that, for any item $e$, $v_{1}(e)\in\{1,\frac{1}{\epsilon}\}$ and $v_{2}(e)\in\{\epsilon^{2},\epsilon\}$, where $10<\epsilon<$ 1. When $t=1$, the first item $e_{1}$ with the value of 1 for agent 1 and $\epsilon$ for agent 2 arrives. If it is allocated to agent 2, no further items arrive. Hence, this allocation is EF1, MMS, and $\epsilon$-max-USW. If not, agent $1$ picks it, and this allocation is EF1, MMS, and max-USW. Then, when $t=2$, the second arriving item $e_{2}$ has the value of 1 for agent 1 and $\epsilon^{2}$ for agent 2. If it is allocated to agent 1, this allocation is 0-EF1, 0-MMS, and max-USW. If it is allocated to agent 2, this allocation is EF1, MMS, and $\frac{\epsilon^{2}+1}{1+\frac{1}{\epsilon}}=\frac{\epsilon^{3}+\epsilon}{1+\epsilon}$-max-USW. For any $\beta>0$, let $\epsilon<\beta^{2}$, we can derive that no online algorithm can guarantee an NW allocation that is $\alpha$-EF1/MMS and $\beta$-max-USW for any $\alpha,\beta>0$
@@ 3.3.2 The Combination with Additive Binary Valuations
In this part, we consider the combination of additive bi-valued and additive binary valuation functions and study the setting where only one agent (agent $n$) has the additive bi-valued valuation function, and the remaining agents have identical additive binary valuation functions. We show that an NW allocation that is EF1 and MMS always exists and can be computed by the Adapted-Picking Algorithm (Algorithm 6). The high level idea of our algorithm is that we allocate the appearing items that all agents value at non-zero values evenly with respect to the requirement that agent $n$ has the top priority of picking the item with the value of $b_{n}$ and the lowest priority of picking the item with the value of $a_{n}$. In that case, agent $n$ takes advantage of the option of items, and then we can guarantee that EF1 and MMS hold simultaneously.
Theorem 7.
For the deterministic allocation of indivisible goods for $(n-1)$ agents with identical additive binary valuation functions and one agent with the additive bi-valued valuation function, the Adapted-Picking Algorithm (Algorithm 6) computes an NW allocation that satisfies EF1 and MMS.
Input: An instance $(T,N,(v_{i})_{i\in N})$ with identical $(n-1)$ additive binary valuation functions $(v_{1},\ldots,v_{n})$ and one additive bi-valued valuation function $(v_{n})$
Output: An non-wasteful EF1 and MMS allocation $A$
@@
2Let $A=(\emptyset,\ldots,\emptyset)$;

### v2

@@ 12    $\alpha=2$;
$\lambda_{2}=2$. For the arriving item $e_{2}$ when $\lambda_{2}=2$, let us consider two cases about $\mu_{2}$.
  • When $\mu_{2}=1$, we need to consider two subcases. If $v_{i}(e_{2})=b_{i}$, agent $i$ picks it. For agent $i$, her new bundle is EF since we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}$ and $v_{i}(e_{2})=b_{i}=v_{i}(e_{1})$. For agent $j$, her bundle is $\frac{1}{2}$-EF1 since we have $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\setminus\{g\})$ for some $g\in A_{j}$ and $v_{j}(e_{1})=b_{j}\geq v_{j}(e_{2})$. Thus, this allocation is $\frac{1}{2}$-EF1 and satisfies the properties in Observation 3. Additionally, we set $\lambda_{2}=1$, $\mu_{2}=0$, and keep $\alpha=3$, implying that the DBC algorithm continues to be executed. If $v_{i}(e_{2})=a_{i}$, it is allocated to agent $j$. For agent $i$, her bundle is EF1 since after removing $e_{1}$ from $A_{j}$,we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}=v_{i}(A_{j}\cup\{e_{2}\})$. For agent $j$, her new bundle is trivially EF since we have $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$, $v_{j}(e_{1})=b_{j}\geq v_{j}(g_{2})$ and $v_{j}(e_{2})\geq a_{i}=v_{j}(g_{1})$. Thus, this allocation is EF1 without an envy cycle, and we set $\alpha=1$, $\lambda_{2}=1$, and $\mu_{2}=0$, implying the EGP algorithm is ready to be executed.
  • When $\mu=2$, it is directly allocated to agent $j$. For agent $i$, her bundle is EF since $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}$ and $v_{i}(e_{1})=b_{i}\geq v_{i}(e_{2})$. For agent $j$, her new bundle is $\frac{1}{2}$-EF1 since $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\setminus\{g\})$ for some $g\in A_{i}$ and $v_{j}(e_{2})\geq v_{j}(e_{1})=a_{j}$. Thus, this allocation is $\frac{1}{2}$-EF1 and satisfies the properties in Observation 3. Additionally, we set $\lambda_{2}=1$, $\mu_{2}=0$, and keep $\alpha=3$, implying that the DBC algorithm continues to be executed.
@@
Input: $N,A,\alpha,e,(\lambda_{1},\mu_{1})$
Output: An NW allocation A, $\alpha$, and $(\lambda_{1},\mu_{1})$
@@
2if $\lambda_{1}=1$ then
3
@@
5
6    $A_{j}=A_{j}\cup\{e\}$;
7    $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$;
8 else
9
10    if $v_{j}(e)=b_{j}$ then
11
12       $A_{j}=A_{j}\cup\{e\}$;
13       $\lambda_{1}=2$ and $\mu_{1}=1$;
14    else
15       $A_{i}=A_{i}\cup\{e\}$;
16       $\lambda_{1}=2$ and $\mu_{1}=2$;
17      end if
18
19   end if
20
21else
22
23 if $\mu_{1}=1$  then
24
25    $A_{i}=A_{i}\cup\{e\}$;
26 else
27
28    if $v_{i}(e)=a_{i}$ then
29
30       $A_{j}=A_{j}\cup\{e\}$;
31    else
32
33       $A_{i}=A_{i}\cup\{e\}$;
34       $\alpha=3$, $\lambda_{1}=1$, and $\mu_{1}=0$;
35      end if
36
37   end if
38
39 $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$;
40 end if
41
return A, $\alpha$ and $(\lambda_{1},\mu_{1})$;
Algorithm 5 PBC (Preliminary-Breaking-Cycle)
Input: $N,A,\alpha,e,(\lambda_{2},\mu_{2})$
Output: An NW allocation A, $\alpha$, and $(\lambda_{2},\mu_{2})$
1
2if $\lambda_{2}\mod 2=1$ then
3
4 if $v_{i}(e)=a_{i}$ then
5    $A_{j}=A_{j}\cup\{e\}$;
6    $\alpha=1$, $\lambda_{2}=1$, and $\mu_{2}=0$;
7 else if $v_{i}(e)=b_{i}$ and $v_{j}(e)=b_{j}$ then
@@ 9    $A_{j}=A_{j}\cup\{e\}$;
10    $\lambda_{2}=\lambda_{2}+1$ and $\mu_{2}=1$;
@@
12    $A_{i}=A_{i}\cup\{e\}$;
13    $\lambda_{2}=\lambda_{2}+1$ and $\mu_{2}=2$;
14   end if
15
16 else if $\lambda_{2}\mod 2=0$ then
17 if  $\mu_{2}=1$  then
18    if $v_{i}(e)=b_{i}$ then
19
20       $A_{i}=A_{i}\cup\{e\}$;
21       $\lambda_{2}=\lambda_{2}+1$ and $\mu_{2}=0$;
@@ 29    $A_{j}=A_{j}\cup\{e\}$;
30    $\lambda_{2}=\lambda_{2}+1$ and $\mu_{2}=0$;
@@
34return A, $\alpha$, $(\lambda_{2},\mu_{2})$;
Algorithm 6 DBC (Deep-Breaking-Cycle)
Proof of Theorem 4. In the AEGP, it can be seen that for different values of the algorithm indicator $\alpha$, we need to choose the corresponding algorithm, i.e., $\alpha=1$: EGP, $\alpha=2$: PBC, or $\alpha=3$: DBC, to execute. By Lemmas 3 and 4, it can be seen that the output allocation is NW and $\frac{1}{2}$-EF1. Regarding MMS, by the implication between EF1 and MMS Amanatidis et al. (2018), i.e., $\alpha$-EF1 $\Rightarrow$ $\frac{\alpha}{(n-1)\alpha+1}$-MMS, where $n$ is the number of agents, it is clear that the output allocation is $\frac{\frac{1}{2}}{\frac{1}{2}+1}=\frac{1}{3}$-MMS.∎
For an independent interest, in the Appendix, we show that if each online item has a deadline of one period, i.e., each arriving item can wait for at most one round to be allocated, EF1 allocation can be found in the above setting. At last, we present the impossible results to show that our approximation guarantees for EF1 and MMS are tight, and the approximate EF1 (or MMS) is incompatible with USW.
@@
For two agents with additive personalized bi-valued valuation functions, no deterministic online algorithm can compute an NW allocation that is $(\frac{1}{2}+\epsilon)$-EF1 or $(\frac{1}{3}+\epsilon)$-MMS for any $\epsilon>0$.
@@
$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$
agent 11 $\epsilon$ 11 $\epsilon$ 1
@@
$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$
agent 11 $\epsilon$ 11 $\epsilon$ $\epsilon$ 11
agent 2 $\epsilon$ $\epsilon$ $\epsilon$ 1 11
@@
$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$
agent 11 $\epsilon$ 11 $\epsilon$ $\epsilon$
@@
Table 6: The impossibility results of EF1 and MMS for additive personalized bi-valued valuations for two agents.
Let $a_{1}=a_{2}=\epsilon$ and $b_{1}=b_{2}=1$, where $0<\epsilon<1$. Consider the following online instance for $n=2$, where the value of each item is shown in Table 6. When $t=1$, the item with the value of $\epsilon$ for both agents arrives. Without loss of generality, agent 1 picks it. When $t=2$, the arriving item has the value of 1 for agent 1 and $\epsilon$ for agent 2. If it is allocated to agent 1, agent 2’s bundle is empty, resulting in 0-EF1 and 0-MMS. Thus, it should be allocated to agent 2. Then, when $t=3$, the arriving item has the value of $\epsilon$ for both agents. Let us consider two different choices of this item.
@@
Given any $\alpha,\beta>0$, for the deterministic allocation of indivisible goods, no online algorithm can compute an NW allocation satisfying
  • $\alpha$-EF1 and $\beta$-max-USW,
  • or $\alpha$-MMS and $\beta$-max-USW,
even for two agents with additive personalized bi-valued valuations.
@@
Without loss of generality, we assume that $v_{i}\in\{1,\frac{1}{\epsilon}\}$ and $v_{2}\in\{\epsilon^{2},\epsilon\}$, where $0<\epsilon<1$. When $t=1$, the first item $e_{1}$ that has the value of 1 for agent 1 and $\epsilon$ for agent 2 arrives. If it is allocated to agent 2, no further items arrive. Hence, this allocation is clearly EF1, MMS, and $\epsilon$-max-USW. If not, agent $1$ picks it, and this allocation is EF1, MMS, and max-USW. Then, when $t=2$, the second arriving item $e_{2}$ has the value of 1 for agent 1 and $\epsilon^{2}$ for agent 2. If it is allocated to agent 1, this allocation is 0-EF1, 0-MMS, and max-USW. If it is allocated to agent 2, this allocation is EF1, MMS, and $\frac{\epsilon^{2}+1}{1+\frac{1}{\epsilon}}=\frac{\epsilon^{3}+\epsilon}{1+\epsilon}$-max-USW. For any $\beta>0$, let $\epsilon<\beta^{2}$, we can derive that no online algorithm can guarantee an NW allocation that is $\alpha$-EF1 (or $\alpha$-MMS) and $\beta$-max-USW for any $\alpha,\beta>0$
@@ 3.3.2 The Combination with Additive Binary Valuations
In this part, we focus on the setting where only one agent (agent $n$) has an additive personalized bi-valued valuation function, while all other agents share identical additive binary valuation functions, formed by $(T,N,(v_{i})_{i\in N\setminus\{n\}},v_{n})$. We show that an NW allocation that is EF1 and MMS always exists and can be computed by the Adapted-Picking Algorithm (Algorithm 7). The high-level idea of our algorithm is to allocate the arriving items that all agents value at non-zero values “evenly”, according to the requirement that agent $n$ has the top priority of picking the item with the value of $b_{n}$ and the lowest priority of picking the item with the value of $a_{n}$. In that case, agent $n$ takes advantage of the item options, and then we can guarantee that EF1 and MMS hold simultaneously. In Table 7, we illustrate how our algorithm works for two agents.
Input: An instance $(T,N,(v_{i})_{i\in N\setminus\{n\}},v_{n})$ with identical (n1)(n-1) additive binary valuation functions $(v_{1},\ldots,v_{n})$ and one additive personalized bi-valued valuation function $v_{n}$
Output: An NW, EF1, and MMS allocation $\bf A$}
@@
2Let ${\bf A}=(\emptyset,\ldots,\emptyset)$;